# HiPreNets: High-Precision Neural Networks through Progressive Training

### v2

@@ 7 Improving Boundary Accuracy through Domain Expansion
We evaluate this approach on Function I.13.12, where 1,000,000 validation samples are generated with all input variables in [1.1,4.9][1.1,4.9][ 1.1 , 4.9 ]. Two residual ensembles are trained: one on 1,000,000 samples from the domain [1.1,4.9][1.1,4.9][ 1.1 , 4.9 ] (the baseline), and another on the slightly expanded domain [1,5][1,5][ 1 , 5 ]. We compare each ensemble’s performance on the validation data in Table 9 and in Figure 12, where we see that the ensemble trained on an expanded domain sees a substantial reduction in RMSE and $L^{\infty}$ norm error.
Function Training Domain
Final RMSE
on [1.1,4.9]
Final $L^{\infty}$ Norm Error
on [1.1,4.9]
@@
Table 9: The final validation RMSE and $L^{\infty}$ norm error for evaluating on data in [1.1,4.9] for one network trained on data in [1,5] and another trained on data in [1.1,4.9]. Averaged across three trials.
Figure 12: Validation results on data generated in [1.1,4.9][1.1,4.9][ 1.1 , 4.9 ] at each iteration. The normal domain network is trained on data generated in [1.1,4.9][1.1,4.9][ 1.1 , 4.9 ], while the expanded domain network is trained on data generated in [1,5][1,5][ 1 , 5 ].
To explore the limits of this approach, we conduct a second experiment with a more pronounced domain expansion. Now, the 1,000,000 validation data samples are all generated to have input variables in [2,4][2,4][ 2 , 4 ]. Again, two residual ensembles are trained with 1,000,000 samples: one on [2,4][2,4][ 2 , 4 ] and the other on the same expanded domain [1,5][1,5][ 1 , 5 ]. The final RMSE and $L^{\infty}$ norm error for each ensemble is given in Table 10. We now observe that the expanded domain ensemble performs noticeably worse in RMSE than the original domain ensemble. This suggests that when the expanded region is too large relative to the target domain, the model may underfit the region of interest unless given a proportionally larger training set.
Function Training Domain
Final RMSE
on [2,4]
Final $L^{\infty}$ Norm Error
on [2,4]
@@
Table 10: The final validation RMSE and $L^{\infty}$ norm error for evaluating on data in [2,4] for one network trained on data in [1,5] and another trained on data in [2,4]. Averaged over three trials.
Figure 13: Validation results on data generated in [2,4][2,4][ 2 , 4 ] at each iteration. The normal domain network is trained on data generated in [2,4][2,4][ 2 , 4 ], while the expanded domain network is trained on data generated in [1,5][1,5][ 1 , 5 ].
In summary, expanding the training domain can significantly reduce boundary errors when the expanded region is modest and representative. However, overly broad expansion may lead to underfitting in the target domain unless training density is increased or weighted sampling is applied. This approach could potentially be expanded to applications where it is not possible to collect data outside the target domain. For example, in climate prediction models, it may be beneficial to extrapolate existing data beyond its limits to include some “unfeasible” cases.
@@ 8 Conclusion
Training highly accurate neural networks for scientific and engineering tasks remains challenging due to the complex interplay of non-convex optimization, sensitivity to hyperparameters, and the lack of a structured framework for improving worst-case error. Inspired by the principles of gradient boosting from classical machine learning, our approach decomposes the problem of fitting a complex function into a sequence of simpler tasks, where each stage of training focuses on correcting the residual errors of its predecessor. This method not only reduces the dependence on extensive hyperparameter tuning, but also offers a modular and more scalable path toward model refinement.
Unlike traditional approaches that rely heavily on minimizing average error via mean squared loss, our method also prioritizes control over the $L^{\infty}$ norm error, which is crucial in domains that require safety-critical decision-making. Through extensive experiments on regression problems drawn from the Feynman dataset, we demonstrated that our HiPreNet framework consistently improves both RMSE and $L^{\infty}$ norm error, outperforming standard fully connected networks and even more advanced architectures like Kolmogorov–Arnold Networks (KANs).
We further explored design strategies that enhance this framework to further reduce both the RMSE and $L^{\infty}$ norm error, such as adaptively increasing network width across residual stages, using loss functions targeted at maximum error reduction, and applying targeted techniques such as local patching and training domain expansion to focus on correcting regions of greatest error. These strategies do not always offer guaranteed improvement, but offer additional ways to approach the design of a HiPreNet framework that may be suitable for a given task.
In future work, it would be worthwhile to explore if there are certain properties of the residuals that can be taken advantage of to determine the optimal size of each refinement network. It would additionally be worthwhile to modify the framework to handle tasks such as time series data or classification. The HiPreNet framework can also be extended to more complex architectures such as convolutional neural networks or physics informed neural networks, enabling improved performance on computer vision problems or data governed by physical laws.

### v3

@@ 5 Improving Boundary Accuracy through Domain Expansion
We evaluate this approach on Function I.13.12, where 1,000,000 validation samples are generated with all input variables in $[1.1,4.9]$. Two residual ensembles are trained: one on 1,000,000 samples from the domain $[1.1,4.9]$ (the baseline), and another on the slightly expanded domain $[1,5]$. We compare each ensemble’s performance on the validation data in Table 13 and in Figure 13, where we see that the ensemble trained on an expanded domain sees a substantial reduction in RMSE and $L^{\infty}$ norm error.
Function Training Domain Final RMSE on [1.1,4.9] Final $L^{\infty}$ Norm Error on [1.1,4.9]
@@
Table 13: Validation RMSE and $L^{\infty}$ norm error on $[1.1,4.9]$ for networks trained on the same domain versus a modestly expanded domain $[1,5]$. Slight expansion reduces both error metrics.
Figure 13: Residuals at each HiPreNet stage for Function I.13.12, evaluated on validation data in $[1.1,4.9]$. The standard network (left column) is trained on the same domain, while the expanded network (right column) is trained on the larger domain $[1,5]$. The expanded training domain produces smaller, more uniform residuals by reducing boundary errors.
To explore the limits of this approach, we conduct a second experiment with a more pronounced domain expansion. Now, the 1,000,000 validation data samples are all generated to have input variables in $[2,4]$. Again, two residual ensembles are trained with 1,000,000 samples: one on $[2,4]$ and the other on the same expanded domain $[1,5]$. The final RMSE and $L^{\infty}$ norm error for each ensemble is given in Table 14. We now observe that the expanded domain ensemble performs noticeably worse in RMSE than the original domain ensemble. This suggests that expanding the domain dilutes training density per unit volume in the target region. For a fixed training set size, the expanded model must distribute capacity over a larger input space. Therefore, when the expanded region is too large relative to the target domain, the model may underfit the region of interest unless given a proportionally larger training set.
Function Training Domain Final RMSE on [2,4] Final $L^{\infty}$ Norm Error on [2,4]
@@
Table 14: Validation RMSE and $L^{\infty}$ norm error on $[2,4]$ comparing training on the exact target domain versus a much larger domain $[1,5]$. Overly broad expansion degrades both error metrics due to diluted training density.
Figure 14: Residuals at each HiPreNet stage for Function I.13.12, evaluated on validation data in $[2,4]$. The standard network is trained on the same domain, while the expanded network is trained on the much larger domain $[1,5]$. The expanded domain here performs worse, as the large gap between training and validation domains dilutes training density and degrades accuracy in the target region.
In summary, expanding the training domain can significantly reduce boundary errors when the expanded region is modest and representative. However, overly broad expansion may lead to underfitting in the target domain unless training density is increased proportionally. Therefore, this technique is most naturally applicable in settings where data generation or collection is cheap and where there is a well-defined boundary of the domain of interest so that a modest expansion can be correctly designed.
6 High Dimensional ODE Example
We next evaluate the HiPreNet framework on a higher dimensional problem: a 10-generator 39-bus power system which is a reduced-order model of the New England power grid [4]. The dynamics are described by the 20-dimensional dynamical system
$\frac{d\omega_{i}}{dt}$ $=\frac{\omega_{0}}{2H_{i}}\Bigl(P_{m}-D\frac{\omega_{i}-\omega_{0}}{\omega_{0}}-E_{i}^{2}G_{ii}$
$\quad-\sum_{j=1,j\neq i}^{10}E_{i}E_{j}\left[B_{ij}\sin(\delta_{i}-\delta_{j})+G_{ij}\cos(\delta_{i}-\delta_{j})\right]\Bigr)$
$\frac{d\delta_{i}}{dt}$ $=\omega_{i}-\omega_{0}$
for $i=1,\dots,10$, where $\omega_{i}$ is the rotor speed in rad/s, and $\delta_{i}$ is the rotor angle in rad. The other parameters are the inertial constant of the generator $H_{i}$, the synchronous angular frequency $\omega_{0}=2\pi f_{0}$ in rad/s for an ac power system with frequency $f_{0}$, the damping coefficient $D$, the mechanical power input from the turbine $P_{m}$, and the electromotive force or internal voltage of the generator $E_{i}$. In addition, the mutual admittance $G_{ij}+jB_{ij}$ between $E_{i}$ and $E_{j}$ is the $i$th row and $j$th column element of the admittance matrix among all electromotive forces, and $G_{ii}$ is the conductance representing the local load seen from $E_{i}$.
We aim to learn a flow map that predicts the state $x(t+5)$ directly from $x(t)$. Neural networks are well suited for this task, as they can approximate high dimensional nonlinear state-to-state mappings and capture the effects of nonlinear interactions over finite time horizons. The five-second prediction horizon makes this task challenging, as the learned flow map must encode long-horizon nonlinear dynamics. High accuracy is therefore essential, since even modest approximation errors may be unacceptable when the flow map is highly nonlinear.
There are several motivating reasons for this choice of a coarse prediction horizon. First, many power system analysis tasks require the system state after several seconds rather than at every smaller interval. Such a task might be assessing whether the system has returned to a near-steady operating point following a disturbance. A 5 second neural network surrogate directly outputs this quantity in a single forward pass, without the storage or processing of intermediate states that a finer resolution surrogate would require, which enables fast inference. Additionally, a finer resolution surrogate applied autoregressively would require many sequential network evaluations, each introducing error that compounds over time. A coarser surrogate avoids this accumulation entirely and achieves a large computational speedup precisely because it bypasses many of these intermediate timesteps in a single evaluation.
To generate data, we numerically solve this system using the Forward Euler method for 500 different initial conditions drawn from the uniform distribution $R$ where
$R=$ $\{x\in\mathcal{R}^{20};\ -0.4\pi<\delta_{i}-(\delta_{0})_{i}<0.4\pi$
$|\omega_{i}-120\pi|<1.5\ \text{for}\ 1\leq i\leq 10\}$
using a timestep of $dt=\frac{1}{1000}$ for 20 seconds. From these trajectories, we collect each valid pair of states $\left(x(t),x(t+5)\right)$ for time $t\in[0,15]$, resulting in approximately $7.5$ million samples.
To learn this flow map, we first train a neural network with four hidden layers of 128 neurons each on 80% of the total samples. We then apply a single refinement stage to test whether residual correction improves finite-horizon prediction in high dimensional dynamical systems. This was accomplished by training a refinement network with four hidden layers of 256 neurons on the dataset $\left(x(t),r(t)\right)$ where $r(t)$ is the base network’s residuals standardized using the residuals’ mean and standard deviation. Both networks were optimized using Adam with MSE loss for 20 total epochs. Adam was chosen over BFGS for this problem due to its scalability benefits for higher dimensional systems.
The RMSE and $L^{\infty}$ norm error after evaluating the model on the remaining 20% of samples are shown in Table 15. The refinement network successfully reduces the observed RMSE for $\omega$ by approximately $60\%$ and for $\delta$ by approximately $70\%$, and reduces the $L^{\infty}$ norm error for $\omega$ by approximately $35\%$ and for $\delta$ by approximately $58\%$.
Base $\omega$ Final $\omega$ Base $\delta$ Final $\delta$
RMSE (rad/s or rad) 4.130e-03 1.670e-03 5.647e-03 1.484e-03
$L^{\infty}$ Error (rad/s or rad) 2.765e-01 1.802e-01 1.242e-01 5.252e-02
Table 15: Validation RMSE and $L^{\infty}$ norm error for rotor speed $\omega$ and rotor angle $\delta$ for the base and refined networks. Refinement reduces $\omega$ RMSE by approximately $60\%$, $\delta$ RMSE by approximately $70\%$, $\omega$ $L^{\infty}$ error by approximately $35\%$, and $\delta$ $L^{\infty}$ error by approximately $58\%$.
Additionally, for each validation sample, with error defined as the $L^{\infty}$ norm error across the 10 generators, we report quantiles of this error for the map from $x(t)$ to $x(t+5)$ in Figure 15 for the rotor angle $\delta$ and in Figure 16 for the rotor speed $\omega$. These results indicate that refinement reduces both average error as well as rare, high-magnitude prediction errors for both variables.
Figure 15: Quantiles of the per-sample maximum prediction error across all 10 generators for the rotor angle $\delta$, evaluated on the validation data. Each point represents the error threshold below which that percentage of samples falls. The refined model consistently reduces both typical and worst-case errors relative to the base model.
Figure 16: Quantiles of the per-sample maximum prediction error across all 10 generators for the rotor speed $\omega$, evaluated on the validation data. Each point represents the error threshold below which that percentage of samples falls. The refined model consistently reduces both typical and worst-case errors relative to the base model.
6.1 Targeted Data Augmentation as Weighted Sampling
From analyzing the residuals of the base model, we observe that the largest ones occur for initial states $x(t)$ where $t\in[0,5]$. This is explained by the dynamics of this system being in a transient state for the first few seconds, before settling into a steady state after approximately 10 seconds. These transient dynamics are more difficult to learn than the steady state dynamics, but the vast majority ($\approx 75\%$) of the training dataset is composed of steady state samples. This is similar to the boundary concentration of residuals observed in the low dimensional examples in Section 4.2. However, unlike those cases where the dataset was fixed, here the simulator allows us to generate additional training data cheaply. We therefore apply targeted data generation similar to the approach in Section 4.3.2. For this example, we will generate completely new samples concentrated in the high error transient region.
In practice, we generate approximately 2 million more training samples (which are $(x(t),x(t+5))$ pairs), all of which have $t\in[0,5]$. We then augment the original training dataset with these new samples. This shifts the balance of the dataset to be approximately $40\%$ transient samples and $60\%$ steady state samples.
With this augmented training dataset, we compute the residuals using the base model. The base model remains trained solely on the original dataset. These augmented residuals are now used as the targets for the refinement network, which has the same four hidden layers of 256 neurons architecture, and we again optimize it for 20 total epochs using Adam and MSE loss.
The resulting RMSE and $L^{\infty}$ norm error of the augmentation trained network evaluated on the original base validation dataset are given in Table 16. As shown, the $\omega$ RMSE is reduced by approximately $24\%$ and the $\omega$ $L^{\infty}$ norm error by $51\%$ compared to the results from the network trained solely on the base dataset that are reported in Table 15. Similarly, the $\delta$ RMSE is reduced by approximately $20\%$ and the $\delta$ $L^{\infty}$ norm error by $23\%$ compared to the results from those tables. Overall, these results suggest that augmenting the dataset with more data concentrated around the large residual samples allows for the refinement network to better correct the base network.
Base Dataset $\omega$ Augmented Dataset $\omega$ Base Dataset $\delta$ Augmented Dataset $\delta$
RMSE 1.670e-03 1.276e-03 1.484e-03 1.182e-03
$L^{\infty}$ Error 1.802e-01 8.913e-02 5.252e-02 4.046e-02
Table 16: Validation RMSE and $L^{\infty}$ norm error for rotor speed $\omega$ and rotor angle $\delta$ for the base and refined networks. Refinement reduces $\omega$ RMSE by approximately $24\%$, $\delta$ RMSE by approximately $20\%$, $\omega$ $L^{\infty}$ error by approximately $51\%$, and $\delta$ $L^{\infty}$ error by approximately $23\%$.
6.2 Flow Map Multiple Iterations Forward
We next evaluate how well the trained neural networks can approximate the system state at further times. Given the initial state $x(0)$, we autoregressively apply the trained surrogate to predict $x(t+5)$, $x(t+10)$, $x(t+15)$, and $x(t+20)$, where each prediction is obtained by feeding the previous prediction back into the model as input.
In Figure 17, we plot the RMSE and $L^{\infty}$ norm error for both $\omega$ and $\delta$ as functions of the prediction horizon. Results are shown for the base refinement network as well as the augmented refinement network.
For the rotor speed $\omega$, both the RMSE and $L^{\infty}$ norm error decrease as the prediction horizon increases. Physically, this matches with expected behavior, where the power system is stable and rotor speeds converge to a steady state as the transient dynamics decay. There is less variation in $\omega$ to predict as time progresses, and the trained model correctly captures this.
For the rotor angle $\delta$, the results differ between the two refinement networks. Using the base refinement network, we observe that the $\delta$ RMSE and $L^{\infty}$ norm error increase with the prediction horizon. As $\delta$ is an integrated quantity from $\omega$, this behavior is consistent with integrator drift, where small errors in the predicted $\omega$ accumulate into growing $\delta$ errors over successive model steps.
However, when using the augmented refinement network, we see that the $\delta$ errors remain nearly constant or even decrease with the prediction horizon. One likely factor behind this is that the targeted data augmentation substantially reduced early $\omega$ errors, which limited the effects of error accumulation. This can also be explained by the dynamics of the power system, where trajectories converge to a steady state. When the model accurately captures these dynamics, as the augmented refinement network appears to, this attraction toward the true steady state can overcome the error accumulation from $\omega$ errors, resulting in stable or even decreasing $\delta$ error at longer horizons.
Figure 17: RMSE and $L^{\infty}$ norm error for rotor speed $\omega$ and rotor angle $\delta$ as a function of prediction horizon, comparing the base refinement network against the augmented refinement network. For $\omega$, errors decrease with horizon as the system converges to a steady state. For $\delta$, the base refinement network shows growing errors due to error accumulation, while the augmented refinement network maintains stable or decreasing errors across all horizons.
6.3 Neural Network Surrogate Speedup
By using a neural network surrogate for this model, we also achieve a speedup in wall clock time. We run both the surrogate and Forward Euler solver given 100 different random initial conditions, and compute the mean and standard deviation of the runtime. The surrogate model completes a 5-second prediction in $1.24\pm 0.20$ ms, compared to $294.87\pm 2.30$ ms for the Forward Euler solver using a step size of $dt=\frac{1}{1000}$, representing a $238\times$ reduction in computation time. Similarly, the surrogate model completes a 20-second prediction in $11.01\pm 0.89$ ms, compared to $1033.95\pm 11.51$ ms for the Forward Euler solver using the same step size, which is a $94\times$ reduction in computation time. This lower speedup reflects the four autoregressive model evaluations required for the 20 second prediction. When using a surrogate we do not obtain any of the intermediate states, so this is only beneficial for applications where only the final state matters.
Additionally, the surrogate naturally supports batch inference for running simulations for multiple initial conditions. This results in a massive speedup versus running simulations sequentially, and this speedup can be further improved by moving inference to run on a GPU. In Table 17, we present the mean and per-trajectory runtime across batch sizes of 1, 10, 100 5-second predictions on both a CPU and GPU. The experiments were repeated 50 times. In Table 18, we present similar results for 20-second predictions. In both cases, while the GPU is slower for a batch size of 1, it becomes more efficient in per-trajectory time as the batch size grows.
Device Batch Size Mean $\pm$ Std (ms) Per-Trajectory (ms)
CPU 1 $1.241\pm 0.199$ 1.241
10 $2.112\pm 0.392$ 0.211
100 $13.869\pm 0.256$ 0.139
GPU 1 $3.755\pm 0.392$ 3.755
10 $3.704\pm 0.281$ 0.370
100 $4.204\pm 0.129$ 0.042
Table 17: Surrogate model inference time for a single 5-second prediction step across batch sizes on CPU and GPU. Per-trajectory time is the mean divided by batch size.
Device Batch Size Mean $\pm$ Std (ms) Per-Trajectory (ms)
CPU 1 $1.006\pm 0.892 11$.006
10 $21.033\pm 1.476$ 2.103
100 $61.637\pm 0.559$ 0.616
GPU 1 $13.529\pm 3.909$ 13.529
10 $12.290\pm 1.167$ 1.229
100 $16.933\pm 1.069$ 0.169
Table 18: Surrogate model inference time for a 20-second prediction (4 autoregressive steps) across batch sizes on CPU and GPU. Per-trajectory time is the mean divided by batch size.
@@ 7 Conclusion
Training highly accurate neural networks for scientific and engineering tasks remains challenging due to factors such as non-convex optimization, sensitivity to hyperparameters, and the lack of a structured framework for improving worst-case error. Inspired by the principles of gradient boosting from classical machine learning, our approach decomposes the problem of fitting a complex function into a sequence of simpler tasks, where each stage of training focuses on correcting the residual errors of its predecessor. This method not only reduces the dependence on extensive hyperparameter tuning, but also offers a modular and more scalable path toward model refinement.
Unlike traditional approaches that rely heavily on minimizing average error via mean squared loss, our method also prioritizes control over the $L^{\infty}$ norm error, which is crucial in domains that require safety-critical decision-making. Through extensive experiments on regression problems drawn from the Feynman dataset, we demonstrated that our HiPreNet framework consistently improves both RMSE and $L^{\infty}$ norm error, outperforming standard fully connected networks and even more advanced architectures like Kolmogorov–Arnold Networks (KANs). We additionally showed how the framework can be applied to a higher dimensional power system example, where it allowed for the creation of a fast and accurate neural network surrogate model.
Across both low and high dimensional settings, we observed the consistent theme that the largest errors concentrate in specific, identifiable regions of the input domain. These regions appeared to be near boundaries for the lower dimensional Feynman functions and in the transient regime for the power system model. We addressed this through three related techniques. The first was weighted sampling, which resamples existing data proportionally to residual magnitude. We secondly investigated localized patching as a preliminary technique, which trains a simple neural network to correct a localized spike in the residuals. Thirdly, we tried targeted data augmentation, which generates new samples in the high error region when new data collection is possible. These techniques represent a spectrum of approaches to the same underlying problem, with the appropriate choice depending on whether new data can be generated and whether the high error region has identifiable structure.
In future work, it would be beneficial to explore if there are certain properties of the residuals that can be taken advantage of to determine the optimal size of each refinement network. It would also be worthwhile to modify the framework to handle classification tasks. The HiPreNet framework could also be applied to different architectures such as transformers and neural operators, where investigating whether the residual correction principle remains effective given their different training dynamics appears to be a promising future direction.